OpenStudy (anonymous):

Which system has no solutions? (Points : 4) x > 3 x < 1 x > 3 x > 1 x < 3 x < 1 x < 3 x > 1

OpenStudy (alexandervonhumboldt2):

only firts because there is nomber that d more than 3 but less than one

OpenStudy (alexandervonhumboldt2):

1

OpenStudy (anonymous):

Ok so x>3 x<1

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

ok what about this one. Solve the following system of equations. x + y = 12 x – y = 6 ?

OpenStudy (alexandervonhumboldt2):

x=12-y x=6+y

OpenStudy (alexandervonhumboldt2):

12-y=6+y

OpenStudy (alexandervonhumboldt2):

-2y=-6

OpenStudy (alexandervonhumboldt2):

y=3

OpenStudy (anonymous):

(9,3) ?

OpenStudy (alexandervonhumboldt2):

x=9

OpenStudy (anonymous):

okay! so heres one more.Which equation will result if the substitution method is used to solve the following system of linear equations? x – y = 7 x = y + 3

OpenStudy (anonymous):

????